OpenStudy (anonymous):

what is the value of x in the proportion x+1 over x+3 = 15 over 21?

OpenStudy (mathstudent55):

\(\dfrac{x + 1}{x + 3} = \dfrac{15}{21} \) First, reduce the right fraction. Divide both the numerator and denominator by 3. \(\dfrac{x + 1}{x + 3} = \dfrac{5}{7} \) Now cross multiply. \(5(x + 3) = 7(x + 1) \) Distribute on both sides: \(5x + 15 = 7x + 7\) Subtract 7x from both sides and subtract 15 from both sides. \(-2x = -8\) Divide both sides by -2. \(x = 4\)